40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1214     Entry time: Fri Jan 2 18:49:54 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: LSC modulation frequencies adjusted 
I noticed that the IFO did not lock in the MICH configuration.
This was because AS166Q signal was too small.
The demodulation phase seemed not right, i.e. the I-phase signal was larger than Q.
I suspected that the 166MHz modulation frequency was not exactly on the MC FSR, since I just
recovered the number written on the Marconi after the power failure.
I measured the optimal frequency by the method explained in elog:752.
It was 165981500Hz, which is pretty close to the number Rob measured in elog:952, but significantly different from
the label on the Marconi.
I set the frequencies of all the MARCONIs accordingly and updated the labels.

After this, the AS166 demodulation phase was still not good enough (the Q and I signals were about the same).
So I rotated the phase by 45deg. In principle, this should set the demod-phase right for DARM too. Is it correct, Rob ?
I also adjusted the PD offsets. After those adjustments, MICH locks stably with a slightly increased gain (20 as compared to 10 before).
ELOG V3.1.3-